# Intrinsic Diophantine approximation: a solution to Mahler’s problem

### v1

@@ 2.2.1. Basic definitions and facts
A map $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a contracting similarity if there exist $\lambda\in\mathbb{R}\setminus\left\{0\right\}$ with $|\lambda|<1$, $O\in\mathcal{O}_{d}(\mathbb{R})$ and $a\in\mathbb{R}^{d}$ such that
@@
[…]
@@ 2.2.2. Separation properties of rational IFSs
Although it is not stated explicitly, it is a direct consequence of Theorem 2.2 that if $T$ satisfies the WSC, then $K_{T}$ is Ahlfors regular. For the sake of completeness, we prove this fact in the appendix.
@@
3. Proof of the upper bound in Theorem 1.1
Let $m,b\in\mathbb{N}$, let $p_{1},\dots,p_{m}\in\mathbb{Z}$, and let
@@
Denote by $K_{T}$ the attractor of $T$. In this section, we prove that if $K_{T}$ has empty interior, then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$ such that
$\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q}\leq 1,$
writing
$E_{\psi,T}=\Big\{x\in K_{T}:|x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T}\Big\},$
one has
$\dim_{H}E_{\psi,T}\leq\frac{\dim_{H}K_{T}}{\delta_{\psi}}.$
Together with Theorem 1.4 (see Corollary 1.5), this proves Theorem 1.1.
Let $\widetilde{K}_{T}\subset\mathbb{T}^{1}$ be the projection of $K_{T}$ onto the one-dimensional torus. It is easily verified that $x\in\widetilde{K}_{T}$ implies $bx\in\widetilde{K}_{T}$. Moreover, $\widetilde{K}_{T}$ is compact; hence, if $\widetilde{K}_{T}\neq\mathbb{T}^{1}$, then $\mathbb{T}^{1}\setminus\widetilde{K}_{T}$ is a non-empty open set, while if $\widetilde{K}_{T}=\mathbb{T}^{1}$, then $K_{T}$ has non-empty interior.
We will prove the following result.
@@
Let $b\in\mathbb{N}$ and let $S\subset\mathbb{T}^{1}$ be a $\times b$-invariant set. For every $n\in\mathbb{N}$, set
$Q_{b,n}=\Big\{b^{n-1}<q\leq b^{n}:\ \exists p\in\mathbb{N},\ \frac{p}{q}\in S,\ p\wedge q=1\Big\}.$
Assume that $\mathbb{T}^{1}\setminus S$ has non-empty interior. Then
$\lim_{n\to+\infty}\frac{\log\#Q_{b,n}}{n}=0.$
Before proving Theorem 3.1, we explain how it implies Theorem 1.1.
Let $\psi:\mathbb{N}\to\mathbb{R}_{+}$ be non-increasing. Observe that
$E_{\psi,T}$ $=\Big\{x\in K_{T}:|x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{T},\ p\wedge q=1\Big\}.$
Lemma 3.2.
Let $q\in\mathbb{N}$. Then there exists a constant $C>0$ such that
$\Big\{p\in\mathbb{N}:\ \frac{p}{q}\in K_{T}\Big\}\leq Cq^{\dim_{H}K_{T}}.$
@@
For $p_{1}\neq p_{2}$ with $\frac{p_{1}}{q},\frac{p_{2}}{q}\in K_{T}$, one has
$\Big|\frac{p_{1}}{q}-\frac{p_{2}}{q}\Big|\geq\frac{1}{q},$
so the balls $B(\frac{p_{1}}{q},\frac{1}{2q})$ and $B(\frac{p_{2}}{q},\frac{1}{2q})$ are pairwise disjoint and centered in $K_{T}$. Since, by Proposition 2.3, $K_{T}$ is Ahlfors regular, there exists $\mu\in\mathcal{M}(K_{T})$, a $\dim_{H}K_{T}$-Ahlfors regular probability measure supported on $K_{T}$. Then
$\sum_{p:\ \frac{p}{q}\in K_{T}}\mu\Big(B\Big(\frac{p}{q},\frac{1}{2q}\Big)\Big)\leq 1,$
which implies the desired bound, with $C$ depending only on $\mu$. ∎
We are now ready to establish the upper bound. For any $N\in\mathbb{N}$,
$E_{\psi,T}\subset\bigcup_{n\geq N}\ \bigcup_{\begin{subarray}{c}q\in Q_{b,n}\\ p:\ \frac{p}{q}\in K_{T}\end{subarray}}B\Big(\frac{p}{q},\psi(b^{n})\Big).$
@@
$s_{\varepsilon}=\frac{\dim_{H}K_{T}+\varepsilon}{\delta_{\psi}}+\varepsilon.$
Then, by Lemma 3.2, there exists $\kappa>0$ such that
$\sum_{q\geq N_{\varepsilon}}\sum_{p:\ \frac{p}{q}\in K_{T}}\psi(q)^{s_{\varepsilon}}$ $\leq\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}+1}\kappa\,b^{n\dim_{H}K_{T}}\,b^{\varepsilon n}\,b^{-n(1-\varepsilon)\delta_{\psi}s_{\varepsilon}}$
$=\kappa\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{-n(\delta_{\psi}s_{\varepsilon}-\dim_{H}K_{T}-\varepsilon)}$
$\leq\kappa\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{-n\varepsilon\delta_{\psi}}<+\infty.$
Hence $\dim_{H}E_{\psi,T}\leq s_{\varepsilon}$, and letting $\varepsilon\to 0$ yields the claim. Before proving Theorem 3.1, we recall basic definitions and results related to discrepancy of sequences on $\mathbb{T}^{1}$.

### v3

@@ 2.2.1. Basic definitions and facts
A map $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a contracting similarity if there exist $\lambda\in(-1,1)\setminus\left\{0\right\}$, $O\in\mathcal{O}_{d}(\mathbb{R})$ and $a\in\mathbb{R}^{d}$ such that
@@
$\begin{cases}\Lambda=\{1,\dots,m\},\\ \Lambda^{*}=\bigcup_{k\geq 1}\Lambda^{k},\\ 0<c_{1},\dots,c_{m}<1\text{ the contraction ratios of }f_{1},\dots,f_{m},\\ \text{for }\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*},\ c_{\underline{i}}=c_{i_{1}}\cdots c_{i_{n}},\\ \text{and }f_{\underline{i}}=f_{i_{1}}\circ\cdots\circ f_{i_{n}}.\end{cases}$
@@ 2.2.2. Separation properties of rational IFSs
For $0\leq s\leq d,$ recall that a measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ is called $s$-Alfhors-regular if there exists a constant $C>0$ such that for every $x\in\mathrm{supp}(\mu)$ and every $r>0,$ one has
$C^{-1}r^{s}\leq\mu\Big(B(x,r)\Big)\leq Cr^{s}.$
Although it is not stated explicitly, it is a direct consequence of Theorem 2.2 that if $T$ satisfies the WSC, then $K_{T}$ is Ahlfors regular (i.e. there exists a $\dim_{H}K_{T}$-Alfors-regular measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ with $\mathrm{supp}(\mu)=K_{T}$). For the sake of completeness, we prove this fact in the appendix.
@@
3. Proof of Theorem 1.1 and Theorem 1.5
Let $m,b\geq 2$, let $p_{1},\dots,p_{m}\in\mathbb{Z}$, and let
@@
Denote by $K_{T}$ the attractor of $T$.
Let $\widetilde{K}_{T}=K_{T}\text{ mod }1\subset\mathbb{T}^{1}$ be the projection of $K_{T}$ onto the one-dimensional torus. It is easily verified that if $x\in\widetilde{K}_{T}$, $bx\in\widetilde{K}_{T}$. Moreover, $\widetilde{K}_{T}$ is compact; hence, if $\widetilde{K}_{T}\neq\mathbb{T}^{1}$, then $\mathbb{T}^{1}\setminus\widetilde{K}_{T}$ is a non-empty open set, while if $\widetilde{K}_{T}=\mathbb{T}^{1}$, then $K_{T}$ has non-empty interior.
We will show the following.
@@
Let $S\subset\mathbb{T}^{1}$ be a $\times b$-invariant set. Assume that $\mathbb{T}^{1}\setminus S$ has non-empty interior. Then, there exists $Q\in\mathbb{N}$ such that for every $q\geq Q,$ $q\wedge b=1,$
$(\exists p\in\mathbb{Z},p\wedge q=1:\ \frac{p}{q}\in S)\Rightarrow O_{q}(b)\leq q^{f(q)},$
with $f(q)=\frac{C}{\sqrt{\log\log\log q}},$ where $C>0$ is a constant.
In the next subsection, we explain how Theorem 1.1 and Theorem 1.5 are derived from Theorem 3.1, Theorem 1.6 and Conjecture 1.4. In Subsection 3.2, we prove Theorem 1.6. The two next subsections are dedicated to the proof of Theorem 3.1. Subsection 3.3 recall basic facts and results about equi-distribution theory and Theorem 3.1 is established in Subsection 3.4.
3.1. Proof of Theorem 1.1 and Theorem 1.5 assuming Theorems 3.1, 1.6 and 1.8
3.1.1. Proof of the upper-bound
Given $S\subset\mathbb{T}^{1}$, a $\times b$ invariant set and $\mathcal{N}$ a subset of $\mathbb{N},$ define
$Q_{b,n,\mathcal{N}}(S)=\Big\{q\in\mathcal{N},b^{n-1}<q\leq b^{n}:\ \exists p\in\mathbb{N},\ \frac{p}{q}\in S,\ p\wedge q=1\Big\}$
and $Q_{b,n}=Q_{b,n,\mathbb{N}}.$
Theorem 3.1 together with Theorem 1.6 and Conjecture 1.4 yields the following corollary.
Corollary 3.2.
Let $S\subset\mathbb{T}^{1}$ be a $\times b$-invariant set. Assume that $\mathbb{T}^{1}\setminus S$ has non-empty interior. Then:
  • (1)
    for every $N\in\mathbb{N},$
    $\lim_{n\to+\infty}\frac{\log\#Q_{b,n,\mathcal{P}_{N}}(S)}{n}=0.$
  • (2)
    assume that Conjecture 1.4 holds true, then
    $\lim_{n\to+\infty}\frac{\log\#Q_{b,n}(S)}{n}=0.$
We start by establishing the following upper-bound
Proposition 3.3.
[…]
$\lim_{n\to+\infty}\frac{\log\#Q_{b,n,\mathcal{N}}}{n}=0.$
Then, for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$,
$\dim_{H}\left\{x\in\mathbb{T}^{1}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{N},p\in\mathbb{Z}:\ \frac{p}{q}\in S\right\}\leq\frac{\overline{\dim}_{B}S}{\max\left\{1,\delta_{\psi}\right\}},$
where $\overline{\dim}_{B}(S)$ denotes the upper box dimension of the set $S$.
The upper-bound for Theorem 1.1 and Theorem 1.5 would then readily follow from Corollary 3.2 with $\mathcal{N}=\mathcal{P}_{N}$ and $\mathcal{N}=\mathbb{N}.$
@@ 3.1.1. Proof of the upper-bound
Without loss of generality, we may assume that $\delta_{\psi}\geq 1.$ Observe that, by monotonicity of $\psi,$
$\left\{x\in\mathbb{T}^{1}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{N},p\in\mathbb{Z}:\ \frac{p}{q}\in S\right\}$
$=$ $\left\{x\in\mathbb{T}^{1}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{N},p\in\mathbb{Z},p\wedge q=1:\ \frac{p}{q}\in S\right\}.$
Fix $\varepsilon>0$ and $r>0$ small enough so that, for every set $\mathcal{C}$ of pairwise disjoint balls centered on $S$ of radii $0<\rho<r,$
$\mathcal{C}\leq\Big(\frac{1}{\rho}\Big)^{\overline{\dim}_{B}S+\varepsilon}.$
As for any, $q\in\mathbb{N},$ for any $p_{1}\neq p_{2}$ such that $\frac{p_{1}}{q},\frac{p_{2}}{q}\in S,$ one has
$B(\frac{p_{1}}{q},\frac{1}{2q})\cap B(\frac{p_{2}}{q},\frac{1}{2q})=\emptyset,$
for every $q>\frac{2}{r},$
$\left\{p\in\mathbb{Z}:\ \frac{p}{q}\in S\right\}\leq(2q)^{\overline{\dim}_{B}S+\varepsilon}$
Remark also that for any $N\in\mathbb{N}$,
$\left\{x\in\mathbb{T}^{1}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{N},p\in\mathbb{Z},p\wedge q=1:\ \frac{p}{q}\in S\right\}$
$\subset\bigcup_{n\geq N}\ \bigcup_{\begin{subarray}{c}q\in Q_{b,n,\mathcal{N}}(S)\\ p:\ \frac{p}{q}\in S\end{subarray}}B\Big(\frac{p}{q},\psi(b^{n})\Big).$
@@
$s_{\varepsilon}=\frac{\overline{\dim}_{B}S+2\varepsilon}{\delta_{\psi}}+\varepsilon.$
Then,
$\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}\sum_{q\in Q_{b,n,\mathcal{N}}(S)}\sum_{p:\ \frac{p}{q}\in S}\psi(q)^{s_{\varepsilon}}$ $\leq 2^{\overline{\dim}_{B}(S)+\varepsilon}\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{n(\overline{\dim}_{B}S+\varepsilon)}\,b^{\varepsilon n}\,b^{-n(1-\varepsilon)\delta_{\psi}s_{\varepsilon}}$
$=2^{\overline{\dim}_{B}(S)+\varepsilon}\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{-n(\delta_{\psi}s_{\varepsilon}-\overline{\dim}_{B}S-2\varepsilon)}$
$\leq 2^{\overline{\dim}_{B}(S)+\varepsilon}\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{-n\varepsilon\delta_{\psi}}<+\infty.$
Hence
$\left\{x\in\mathbb{T}^{1}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{N},p\in\mathbb{Z},p\wedge q=1:\ \frac{p}{q}\in S\right\}\leq s_{\varepsilon}$
and letting $\varepsilon\to 0$ yields the claim.
3.1.2. Proof of the lower-bound
Notice that the lower-bound for Theorem 1.5 is implied by Corollary 1.9. Thus we only prove the lower-bound required to establish Theorem 1.1. We focus first on item $(2)$ of Theorem 1.1. Let us fix $N\geq\omega(b).$ Let $r=\frac{p}{b^{k}}\in K_{T}$ be a $b$-adic number. Remark that, for every $n\in\mathbb{N}$ and every $\underline{i}=(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n},$ there exists $p^{\prime}\in\mathbb{Z}$ such that
$f_{\underline{i}}(r)=\frac{p^{\prime}}{b^{n+k}},$
hence
$\omega(h(f_{\underline{i}}(r)))\leq\omega(b).$
As $N\geq\omega(b),$
$\left\{x\in K_{T}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{P}_{N},p\in\mathbb{Z},p\wedge q=1:\ \frac{p}{q}\in K_{T}\right\}$
$\supset\left\{x\in\mathbb{T}^{1}:\ |x-\frac{p}{b^{t}}|\leq\psi(q)\text{ f.i.m. }t\in\mathbb{N},p\in\mathbb{Z},\ \frac{p}{b^{t}}\in K_{T}\right\}$
$=\limsup_{t\in\mathbb{N},p\in\mathbb{Z},\ \frac{p}{b^{t}}\in K_{T}}B\Big(\frac{p}{b^{t}},\psi(b^{t})\Big).$
Moreover, the argument provided in [12, P.1821] yields
$\dim_{H}\limsup_{n\geq 1,\underline{i}=(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}}B(f_{\underline{i}}(r),\psi(n))\geq\frac{\dim_{H}K_{T}}{\delta_{\psi}}$
so that
$\dim_{H}\left\{x\in K_{T}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }q\in\mathcal{P}_{N},p\in\mathbb{Z},p\wedge q=1:\ \frac{p}{q}\in K_{T}\right\}\geq\frac{\dim_{H}K_{T}}{\delta_{\psi}}.$
We now establish item $(1):$
Without loss of generality, assume that $p_{1}\neq 0$ (otherwise we can apply item $(2)$) and write
$r_{0}=\pi\Big((1)_{\infty}\Big)=\sum_{k\geq 1}p_{1}b^{-k}=\frac{p_{1}}{b-1}\in K_{T}\cap\mathbb{Q}.$
For every $n\in\mathbb{N}$ and every $\underline{i}=(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n},$ there exists $p^{\prime}\in\mathbb{Z}$ such that
$f_{\underline{i}}(r_{0})=\frac{p^{\prime}}{b^{n}(b-1)}.$
Since $\omega(h(f_{\underline{i}}(r_{0})))\leq\omega(b^{n}(b-1))=\omega(b)+\omega(b-1),$ the same argument as above yields the desired conclusion.
In the next section, we establish Theorem 1.6.
3.2. Proof of Theorem 1.6
For $b\geq 2,$ $q\in\mathbb{N}$ with $q\wedge b=1,$ writing $q=\prod_{1\leq i\leq\omega(q)}p_{i}^{\alpha_{i}}$ the decomposition in prime factors of $q$, we set
$\widetilde{O}_{q}(b)=\prod_{1\leq i\leq\omega(q)}O_{p_{i}^{\alpha_{i}}}(b).$
One will show the following.
Theorem 3.4.
Let $\varepsilon>0$ be a real number. For every $x\in\mathbb{N}$ large enough,
$\left\{0\leq q\leq x:\ q\wedge b=1,\ \widetilde{O}_{q}(b)\leq q^{\varepsilon}\right\}\leq x^{2\varepsilon}.$
Prior to proving Theorem 3.4, we explain how it implies Theorem 1.6.
Remark that, for any $q\in\mathbb{N}$, $b\wedge q=1$ one has
$O_{q}(b)\leq\widetilde{O}_{q}(b)\leq O_{q}(b)^{\omega(q)}.$
Let $N\in\mathbb{N}$ be an integer. As $f(q)\to 0,$ for every large enough $q$ such that $\omega(q)\leq N$,
$O_{q}(b)\leq q^{f(q)}\Rightarrow\widetilde{O}_{q}(b)\leq q^{Nf(q)}\leq q^{\varepsilon}.$
Thus, Theorem 3.4 yields
$\limsup_{x\to+\infty}\frac{\#\left\{0\leq q\leq x:\ q\wedge b=1,\ \omega(q)\leq N\ O_{q}(b)\leq q^{\varepsilon}\right\}}{\log x}\leq 2\varepsilon$
and letting $\varepsilon\to 0$ proves Theorem 1.6.
We now prove Theorem 3.4 and we start by a lemma.
Lemma 3.5.
Let $p$ be a prime number, coprime with $b$. Write $k_{p}=O_{p}(b)$ and $u_{p}=v_{p}(b^{k_{p}}-1),$ the valuation of $p$ in $b^{k_{p}}-1.$ Then, for every $t\in\mathbb{N},$
$\begin{cases}O_{p^{t}}(b)=k_{p}\text{ if }t\leq u_{p}\\ O_{p^{t}}(b)=k_{p}p^{t-u_{p}}\text{ if }t\geq u_{p}.\end{cases}$
Proof.
For $t\leq u_{p},$ $p^{t}|b^{k_{p}}-1$ and $O_{p^{t}}(b)\geq k_{p}$ yields
$O_{p^{t}}(b)=k_{p}.$
Let us now fix $t>u_{p}.$ Notice that
$b^{k_{p}}=1+mp^{u_{p}},$
with $m\wedge p=1$ (otherwise $p^{u_{p}+1}|b^{k_{p}}-1$). Hence, there exists $\gamma\in\mathbb{N}$ such that
(8) $b^{pk_{p}}=1+m^{p}p^{u_{p}+1}+\gamma p^{u_{p}+2}.$
This yields $O_{p^{u_{p}+1}}(b)|pk_{p}.$ Since $O_{p^{u_{p}+1}}(b)>k,$ $O_{p^{u_{p}+1}}(b)=pk_{p}.$ Moreover, in (8), $m^{p}\wedge p=1,$ so that the same argument applies to $O_{p^{u_{p}+2}}(b)$ and, recursively, to any $O_{p^{k}}(b)$, $k\geq u_{p}$. ∎
Call $\mathcal{P}^{b}$ the set of prime numbers coprime with $b$. For $t\in\mathbb{R},$ we set, when defined,
$S_{t}=\prod_{p\in\mathcal{P}^{b}}\Big(\sum_{k\geq 0}\frac{1}{O_{p^{k}}(b)^{t}}\Big),$
which develops in
$S_{t}=\sum_{n\geq 1,\ n\wedge b=1}\frac{1}{\widetilde{O}_{n}(b)^{t}}.$
For every $p$ prime, write again $k_{p}=O_{p}(b)$ and $u_{p}=v_{p}(b^{k_{p}}-1).$ Then
$\sum_{i\geq 0}\frac{1}{O_{p^{i}}(b)^{t}}=\sum_{k=0}^{u_{p}}\frac{1}{k_{p}^{t}}+\sum_{i\geq u+1}\frac{1}{(k_{p}p^{i-u})^{t}}\leq\frac{u_{p}}{k_{p}^{t}}\Big(\sum_{k\geq 0}\frac{1}{p^{kt}}\Big).$
As $p^{u_{p}}|b^{k_{p}}-1$, for every $p>b$, $u_{p}\leq k_{p}$. Hence, for any such $p$
$\frac{u_{p}}{k_{p}^{t}}\Big(\sum_{k\geq 0}\frac{1}{p^{kt}}\Big)\leq\sum_{k\geq 0}\frac{1}{p^{kt}}$
so that $S_{t}$ converges for every $t>1$.
Fix $1<t<2$ and assume that, for $x\in\mathbb{N},$
$\left\{1\leq n\leq x:\ n\wedge b=1,\ \widetilde{O}_{n}(b)\leq n^{\varepsilon}\right\}>x^{2\varepsilon}.$
Then
$\sum_{1\leq n\leq x:\ n\wedge b=1}\frac{1}{\widetilde{O}_{n}(b)^{t}}\geq\frac{x^{2\varepsilon}}{x^{t\varepsilon}}=x^{2\varepsilon-t\varepsilon}.$
As $x^{(2-t)\varepsilon}\to+\infty,$ this happens for an at most finite set of $x$.
We now proceed to prove Theorem 3.1. The proof is based on the following observation. Fix $\frac{p}{q}\in S$ with $p\wedge q=1$ and $q\wedge b=1$. On the one hand, since $S$ is $\times b$ invariant, for any $k\in\mathbb{N}$ we have $b^{k}\frac{p}{q}\in S$; on the other hand, the sequence $b^{k}p$ should be equidistributed modulo $q$, unless the multiplicative order of $b$ modulo $q$ is very small (see Theorem 3.8 below). Since the complement of $S$ contains an interval, only those $q$ for which $b$ has small multiplicative order $\text{mod }q$ can satisfy $\frac{p}{q}\in S$.
In the next subsection, we recall basic definitions and tools related to the discrepancy of sequences on $\mathbb{T}^{1}$. We then establish Theorem 3.1.